Solve the Equation for Theta :

sec(theta)-1=(root2 - 1)tan(theta)



Answer Given is 2npi and 2npi + pi/4.

Please give detailed steps if possible. (Best Answers will be Approved)

Thanks a Lot !!

Grade:11

2 Answers

Aman Bansal
12 years ago

Dear Ashray,

 ( 1 -  COS X ) / COS X = ( ROOT 2 -1 ) SIN X / ( COS X)

( 1 - COS X) / SIN X = ROOT 2 -1 

 

SQUARING BOTH SIDES

 

(1-COS X) ^ 2 / [ (1+COS X ) * (1-COS X) ] = 3 - 2 ROOT 2

 

(1-COSX)/1+COSX = 3 -2ROOT 2

 1-COSX = 3 - 2ROOT2 +3COSX -2ROOT2 COS X

2ROOT2 - 2 = COS X * ( 4 - 2 ROOT2 )

COS X = ( ROOT 2 - 1) / ( 2 - ROOT 2) 

              =1/ROOT 2

>>>> X= PI/4

Best Of luck

Plz Approve the answer...!!!!

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and winexciting gifts by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple click here to download the toolbar….
So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert


Mayank Jain
12 years ago

Thanks A Lot !!

Really Liked Your Answer !!